# Bibliography.bib

### v1

@@ 8.2. Contact forms
Gromov [gromov1969stable] proved that if $M$ is an open manifold the complete h-principle holds. It tells us in particular, if we restrict ourselves to the case of $\pi_{0}$-surjectivity, that any formal contact form $(\alpha,\beta)$ is homotopic to a contact form $(\tilde{\alpha},d\tilde{\alpha})$ via formal contact forms in an open manifold. If we compare with Corollary 8.3, we see that jiggling only requires a non-degenerate $1$-form $\alpha$ instead of also a $2$-form $\beta$ such that together they form a formal contact form. However, the output of jiggling is also just a piecewise smooth contact form instead of a smooth one.
For closed manifolds $\pi_{0}$-surjectivity was shown to hold by Martinet [Mart] and Lutz [Lutz], but $\pi_{0}$-injectivity was proven to be false by Bennequin [Be]. Hence for closed manifolds the full h-principle does not hold, unless we restrict ourselves to overtwisted contact structures as shown by Eliashberg [El89].
@@ 8.3.1. General position
Figure 8. Example of a triangulation in general position with respect to the horizontal distribution $\xi$ whose crystalline subdivision is not.
@@ 8.3.2. Jiggling
The collection ${\mathcal{C}}_{\Delta}$ is finite for each $\Delta$, as we can construct it as in the proof of Proposition 3.12: by Lemma 3.4 there exists a finite set of model simplices ${\mathcal{C}}^{\prime}_{\Delta}$ for crystalline subdivisions of $\Delta$. Now given a model simplex $\Delta^{\prime}\in{\mathcal{C}}^{\prime}_{\Delta}$, we apply crystalline subdivision at most once to each of its faces. We then use the barycentric cone off to extend the subdivision of the boundary of $\Delta^{\prime}$ to $\Delta^{\prime}$ itself. We add each of the resulting simplices to ${\mathcal{C}}^{\prime}_{\Delta}$, to define the finite collection ${\mathcal{C}}_{\Delta}$. This motivates the following definition.
@@
Let $M$ and $N$ be smooth manifolds and fix a triangulation $T:K\to M$. We define the general position relation ${\mathcal{R}}_{\operatorname{GenPos}}\subset J^{1}(M,N)$ as follows. A jet $\sigma\in J^{1}(M,N)$ with front projection $(x,y)=\pi_{f}(\sigma)\in M\times N$ belongs to ${\mathcal{R}}_{\operatorname{GenPos}}$ if
@@
Fix $M$, $N$, $K$, and $\xi$ as above, and fix a generalized crystalline subdivision $K^{\prime}$ of $K$. Assume that ${\mathcal{C}}_{\Delta}$ is finite for each top dimensional simplex $\Delta\in K$. Then the following hold:
  • The relation ${\mathcal{R}}_{\operatorname{GenPos}}$ is open and fiberwise dense.
  • If $s^{\prime}:M\rightarrow N$ is a piecewise smooth solution of ${\mathcal{R}}_{\operatorname{GenPos}}$ with respect to $K^{\prime}$, then it is in general position with respect to all generalized crystalline subdivisions $K^{\prime\prime}$ of $K$ that also subdivide $K^{\prime}$.
@@
The following is now a consequence of Corollary 6.3. We point out that Definition 1.2 cannot be applied to obtain the following result, since the relation ${\mathcal{R}}_{\operatorname{GenPos}}$ depends on the model simplices for top-dimensional simplices in $K$, which change throughout the proof of Definition 1.2 due to the coning off of arbitrarily many subdivisions of the boundary.
@@
  • $s^{\prime}$ is in general position with respect to $\xi$,
@@
We have already observed that ${\mathcal{C}}_{\Delta}$ is finite for each top-dimensional simplex $\Delta\in K$ and hence by the first item of Lemma 8.8, the relation ${\mathcal{R}}_{\operatorname{GenPos}}$ is open and fiberwise dense. This allows us to jiggle $(s,T)$ using Corollary 6.3. If we check the proof of Corollary 6.3, we see that it produces a sequence of sections $s^{(k)}$ that are piecewise smooth with respect to a triangulation $T:|K^{(k)}|\to M$, and are a piecewise smooth solution over $C_{k}$, where $\{C_{k}\}_{k\in K}$ is an exhaustion by compacts of $M$. Moreover, the functions $s^{(k)}$ and $s^{(k-1)}$ agree over $C_{k-2}$, and $K^{(k)}$ is a generalized subdivision of $K$ and a subdivision of $K^{(k-1)}$. Hence by the second item of Lemma 8.8, we see that the pair $(s^{\prime},T^{\prime})$ produced by Corollary 6.3 satisfies the current statement. ∎
@@ 8.3.3. Triangulations in general position
The above result can then be applied to the identity map ${\operatorname{id}}:M\to M$, which is the setup for Thurston’s jiggling. The resulting jiggled map will be a piecewise immersion by construction. The fact that it is a homeomorphism follows from degree theory and is immediate if we take a small enough jiggling [lurie2009topics, Lecture 4]. Thus:
@@
  • a subcomplex $K^{\prime}\subset K$ such that $T|_{{\mathcal{O}p}(|K^{\prime}|)}$ is in general position,
@@
  • $T^{\prime}$ is in general position with respect to $\xi$, and

### v2

@@ 8.2. Contact forms
[…]
For closed manifolds $\pi_{0}$-surjectivity was shown to hold by Martinet [23] and Lutz [22], but $\pi_{0}$-injectivity was proven to be false by Bennequin [4]. Hence for closed manifolds the full h-principle does not hold, unless we restrict ourselves to overtwisted contact structures as shown by Eliashberg [12].
@@ 8.3.1. General position
Figure 8. Example of a triangulation in general position with respect to the vertical distribution $\xi$ whose subdivision is not.
@@ 8.3.2. Jiggling
The collection ${\mathcal{C}}_{\Delta}$ is finite for each $\Delta$, as we can construct it as in the proof of Proposition 3.14: by Lemma 3.4 there exists a finite set of model simplices ${\mathcal{C}}^{\prime}_{\Delta}$ for crystalline subdivisions of $\Delta$. Now given a model simplex $\Delta^{\prime}\in{\mathcal{C}}^{\prime}_{\Delta}$, we apply crystalline subdivision at most once to each of its faces. We then use the barycentric cone off to extend the subdivision of the boundary of $\Delta^{\prime}$ to $\Delta^{\prime}$ itself. We add each of the resulting simplices to ${\mathcal{C}}^{\prime}_{\Delta}$, to define the finite collection ${\mathcal{C}}_{\Delta}$. This motivates the following definition, of a notion stronger than general position.
@@
Let $M$ and $N$ be smooth manifolds and fix a triangulation $T:K\to M$. We define the very general position relation ${\mathcal{R}}_{\operatorname{VeryGenPos}}\subset J^{1}(M,N)$ as follows. A jet $\sigma\in J^{1}(M,N)$ with front projection $(x,y)=\pi_{f}(\sigma)\in M\times N$ belongs to ${\mathcal{R}}_{\operatorname{VeryGenPos}}$ if
@@
Fix $M$, NN, $K$, and $\xi$ as above, and fix a generalized crystalline subdivision $K^{\prime}$ of $K$. Assume that ${\mathcal{C}}_{\Delta}$ is finite for each top dimensional simplex $\Delta\in K$. Then the following hold:
  • The relation ${\mathcal{R}}_{\operatorname{VeryGenPos}}$ is open and fiberwise dense.
  • If $s^{\prime}:M\rightarrow N$ is a piecewise smooth solution of ${\mathcal{R}}_{\operatorname{VeryGenPos}}$ with respect to $K^{\prime}$, then it is in general position with respect to all generalized crystalline subdivisions $K^{\prime\prime}$ of $K$ that also subdivide $K^{\prime}$.
@@
The following is now a consequence of Corollary 6.3. We point out that Definition 1.2 cannot be applied to obtain the following result, since the relation ${\mathcal{R}}_{\operatorname{VeryGenPos}}$ depends on the model simplices for top-dimensional simplices in $K$, which change throughout the proof of Definition 1.2 due to the coning off of arbitrarily many subdivisions of the boundary.
@@
  • a subcomplex $Q\subset M$ of $T$ such that $s|_{{\mathcal{O}p}(Q)}$ is a solution of ${\mathcal{R}}_{\operatorname{VeryGenPos}}$,
@@
  • $s^{\prime}$ is in (very) general position with respect to $\xi$,
@@
We have already observed that ${\mathcal{C}}_{\Delta}$ is finite for each top-dimensional simplex $\Delta\in K$ and hence by the first item of Lemma 8.8, the relation ${\mathcal{R}}_{\operatorname{VeryGenPos}}$ is open and fiberwise dense. This allows us to jiggle $(s,T)$ using Corollary 6.3. If we check the proof of Corollary 6.3, we see that it produces a sequence of sections $s^{(k)}$ that are piecewise smooth with respect to a triangulation $T:|K^{(k)}|\to M$, and are a piecewise smooth solution over $C_{k}$, where $\{C_{k}\}_{k\in K}$ is an exhaustion by compacts of $M$. Moreover, the functions $s^{(k)}$ and $s^{(k-1)}$ agree over $C_{k-2}$, and $K^{(k)}$ is a generalized subdivision of $K$ and a subdivision of $K^{(k-1)}$. Hence by the second item of Lemma 8.8, we see that the pair $(s^{\prime},T^{\prime})$ produced by Corollary 6.3 satisfies the current statement. ∎
@@ 8.3.3. Triangulations in general position
The above result can then be applied to the identity map ${\operatorname{id}}:M\to M$, which is the setup for Thurston’s jiggling. The resulting jiggled map will be a piecewise immersion by construction. The fact that it is a homeomorphism follows from degree theory and is immediate if we take a small enough jiggling [21, Lecture 4]. Thus:
@@
  • a subcomplex $K^{\prime}\subset K$ such that $T|_{{\mathcal{O}p}(|K^{\prime}|)}$ is in very general position,
@@
  • $T^{\prime}$ is in (very) general position with respect to $\xi$, and